If you have ever driven a sports car and a minivan you know that they don't accelerate the same. What force would it take to accelerate a 609.68 kg sports car from 0 to 60 mph in 4.75 seconds?
Answer #1

Using 1st kinematic equation:

V = U + a*t

a = acceleration of sports car = (V - U)/t

U = initial speed = 0 m/sec

V = final speed = 60 mph = 60 miles/hr = (60 miles/hr)*(1609.34 m/1 miles)*(1 hr/3600 sec) = 26.82 m/sec

t = time taken = 4.75 sec

So,

a = (26.82 - 0)/4.75

a = 5.65 m/sec^2

Now from Newton's 2nd law:

Fnet = m*a

m = mass of sports car = 609.68 kg

Fnet = net force required to accelerate = 609.68*5.65

Fnet = 3444.7 N

Please Upvote.